Solve the exponential equation. Use a calculator to obtain a decimal approximation,correct to two decimal places, for the solution.
4^x+6=6

Answer by stanbon(57361)

Solve the exponential equation. Use a calculator to obtain a decimal approximation,correct to two decimal places, for the solution.
4^(x+6) = 6
---
Take the log of both sides:
(x+6)*log(4) = log(6)
x+6 = [log(6)/log(4)]
x = [log(6)/log(4)]-6
x = -4.708
Cheers,
Stan H.